i am going to a job that uses simple algebra deal in conveting factions of inches to decemial points where can i go to get help on that
Answer by tutorcecilia(2152)

Fractions to decimals
To convert fractions to decimals, divide the numerator by the denominator.
Example:
Convert 1/2 in. to a decimal
1 divided by 2 = 0.5 inches
.
Example: convert 3/7 in. to a decimal
3 divided by 7= .42857 inches
.
Example: convert 7/2 in. to a decimal
7 divided by 2=3.5 in.
.
Convert 3 1/4 in to a decimal:
First re-write the mixed fraction as an improproper fraction:
13/4
13 divided by 4=3.25 in.
.
to convert decimals to a fraction, set-up the decimal as a fraction and reduce.
Example: Convert .25 in to a fraction
.25 is in the 100th place value, so,
25/100 = 1/4 in.
.
Example: to change 3.50 in. to a fraction:
3 is a whole number so set it to the side for now.
.50 is really .5 because the zero doesn't count in this instance.
So, .5 is in the 10th place, 5/10 = 1/2
Putting 1/2 together with the 3 gives 3 1/2
.
Also, try getting a book on elementary math or check the internet under "fractions to decimals".
I hope this helps!